# Stable Trajectory Clustering: An Efficient Split and Merge Algorithm

### v1

@@
Abstract
Clustering algorithms group data points by characteristics to identify patterns. Over the past two decades, researchers have extended these methods to analyze trajectories of humans, animals, and vehicles, studying their behavior and movement across applications. This paper presents whole-trajectory clustering and sub-trajectory clustering algorithms based on DBSCAN line segment clustering, which encompasses two key events: split and merge of line segments. The events are employed by object movement history and the average Euclidean distance between line segments. In this framework, whole-trajectory clustering considers entire entities’ trajectories, whereas sub-trajectory clustering employs a sliding window model to identify similar sub-trajectories. Many existing trajectory clustering algorithms respond to temporary anomalies in data by splitting trajectories, which often obscures otherwise consistent clustering patterns and leads to less reliable insights. We introduce the stable trajectory clustering algorithm, which leverages the mean absolute deviation concept to demonstrate that selective omission of transient deviations not only preserves the integrity of clusters but also improves their stability and interpretability. We run all proposed algorithms on real trajectory datasets to illustrate their effectiveness and sensitivity to parameter variations.
Index Terms:
@@ Abstract
I Introduction
@@
Trajectory clustering is one of the most important data analysis methods developed for analyzing moving objects like humans, animals, and vehicles in various fields and applications. It can be used to identify common behaviors and recognize similar movement patterns, as well as to detect irregular or anomalous movements in different types of trajectory data. Among the use cases of trajectory clustering, we can mention exploring urban travel patterns that reflect the regularity of mobility patterns, which is an important factor in evaluating the optimality of city spatial structure and road network connectivity [1], classifying animal behaviors (foraging, resting, walking) using high-frequency tracking data to study their environmental interactions [2] and clustering animals by species based on their daily movement patterns, which is complicated by the erratic nature of these movements [3]. Trajectory clustering is also used in air traffic categorization [4, 5], traffic network identification [6] and, weather prediction [7].
@@
Anomalies frequently occur in trajectory clustering due to irregularities or deviations in movement patterns. To address this, numerous algorithms have been developed, each designed to effectively detect and analyze these anomalies [8, 9]. Missing, however, is an algorithm that can filter out and ignore insignificant anomalies in trajectories to reveal clustering patterns that are otherwise stable. Researchers have focused in recent years on stable algorithms for clustering data points that select the optimal number of clusters [10]. Said another way, current stable clustering algorithms try to split and merge data points correctly in a way that is robust against random oscillations. Inspired by these developments for point clustering, we propose a stable algorithm for clustering trajectories based on the results of line segment clustering in each time interval. For line segment clustering in a time interval, we use the results from the previous time interval, which reflect the states of moving objects and their changes caused by split and merge events in each time interval. We also present an innovative approach that detects which trajectories can be assigned to their respective cluster despite anomalies, which is a prominent issue in real trajectory data that previous algorithms have not addressed.
@@
Building upon existing trajectory clustering algorithms, we propose algorithms for both whole-trajectory clustering and sub-trajectory clustering. A trajectory can be described as a piecewise linear function, where each segment represents an object’s movement in a specific time interval. Consequently, for trajectory clustering, a set of line segments exists for each time interval, and these sets need to be clustered. For this purpose, Density-Based Spatial Clustering of Applications with Noise (DBSCAN) line segment clustering defines the split and merge procedures of line segments. One key advantage of this approach is that the results from clustering in each time interval can be reused in the subsequent interval, as line segments often retain their previous cluster assignment. In this context, we use the continuous average Euclidean distance to measure the distance between segments. As a result, a data analyst can identify the behaviors of moving objects across all time intervals, determining which objects move together and belong to the same cluster in each interval. The sub-trajectory clustering algorithm focuses on specific time intervals to identify similar sub-trajectories. We employ a sliding window model that utilizes two parameters to recognize temporal patterns and similar sub-sequences. We then present an innovative definition and algorithm for the stable trajectory clustering algorithm, which is a first-of-its-kind advancement. This approach deals with insignificant anomalies and deviations in object movements that obscure broader clustering patterns and can thus be ignored when the goal is quantifying persistent trajectory clusters. It is often important to filter out anomalies in data where imprecision is small relative to the scale of the movement. For instance, in urban traffic analysis, a vehicle making a brief stop at a traffic light or a pedestrian crossing should still be classified within the main traffic cluster. Furthermore, a person briefly deviating from a common trajectory to look at a shop window should not be marked as an outlier in pedestrian movement studies. A related example can be seen in animal movement studies, where a migrating bird or herd of animals temporarily deviates from its migration trajectory to find food or rest, but then returns to their original trajectory afterward. In this algorithm, we employ the mean absolute deviation concept to assess how much each outlier trajectory deviates from its assigned cluster, enabling us to distinguish significant outliers from temporary anomalies. Finally, we apply all algorithms and concepts to two real trajectory datasets and demonstrate the parameter sensitivity and effectiveness of the stable trajectory clustering algorithm based on its impact on data, specifically on how the number of clusters changes after the algorithm is executed. In summary, our contributions are illustrated as a graphical abstract in Fig. 1.
@@
Figure 1: Graphical abstract of the proposed approaches.
II Related Work
@@
In the last two decades, many data point clustering algorithms have been discussed in various fields and applications such as wireless sensor networks [11], flight anomaly detection [12], and image processing [13]. These algorithms try to group similar data points based on their similarities or dissimilarities [14]. There are several types of methods to cluster data points, such as partitioning methods [15, 16, 17], hierarchical methods [18, 19], grid-based methods [20, 21], and density-based methods [22, 23]. Some of these methods are used for clustering data points as well as for trajectory clustering. For example, DBSCAN [22] is a density-based algorithm that is used for clustering both data points and trajectories. It can identify the shape and size of data, can handle noise and outliers efficiently, and does not require the number of clusters as a user-defined input. However, DBSCAN cannot identify clusters with different densities and is sensitive to the choice of user input parameters for clustering.
@@
Trajectory data are recorded in different formats by various devices such as Global Positioning System (GPS) trackers, Radio Frequency Identification (RFID) systems, and smartphones. Numerous algorithms have been designed to cluster these different types of trajectory data. They also contribute to improving research, innovation, and theoretical understanding in various fields and applications. Yu et al. [24] suggested a density-based clustering algorithm for trajectory data streams. They designed an algorithm to handle moving object trajectory data as an incremental line segment stream and introduced a distance measurement approach to cluster those line segments via DBSCAN line segment clustering at the end of the current time interval. Their algorithm also used a structure for updating line segment clusters online. Da et al. [25] dealt with the problem of identifying sub-trajectory clusters at every time interval. To this aim, they defined a structure called a micro-group to take and maintain small groups of moving objects at each time window and presented an algorithm for maintaining each micro-group incrementally to discover patterns. Thus, they suggested a density-based sub-trajectory cluster identification approach by merging micro-groups. Trajectory clustering can be divided into point-based, sub-trajectory-based, and whole-trajectory-based methods. Point-based trajectory clustering is a method of clustering trajectories by comparing specific locations (points) along each trajectory. Instead of looking at the entire trajectory, it focuses on the individual points where the object was at different times, and groups similar ones together. Li et al. [26] offered a group discovery framework to facilitate the efficient online discovery of moving objects traveling collectively. Their framework employed a sampling-independent approach while also naturally supporting the utilization of approximate trajectories. Jensen et al. [27] investigated effectively managing a collection of constantly moving data points in a two-dimensional Euclidean space to maintain a cluster of dynamic data points. Their proposed approach can cluster moving objects incrementally. Tang et al. [28] introduced the smart-and-close discovery algorithm, which efficiently generates companions from trajectory data. In their approach, the objective of the proposed traveling buddy model was to improve the clustering and intersection processes involved in discovering companions. Li et al. [29] recommended utilizing cluster analysis for mobile objects and introduced the notion of moving micro-clusters, which can effectively identify patterns in the mobility of objects even in massive datasets. Gao et al. [30] addressed the challenges of clustering real-time trajectory data generated by GPS-equipped devices. They proposed a novel framework that dynamically models trajectory streams using a self-adaptive k-segment method, considering both spatial and temporal dimensions. This framework supported distributed indexing and processing to handle large-scale data efficiently. It utilized incremental DBSCAN clustering for real-time processing and integrated a two-layered distributed index to enhance scalability. The study’s key contributions included effective trajectory stream modeling, efficient and incremental clustering algorithms, and scalable distributed processing techniques.
@@
Sub-trajectory clustering involves breaking down a trajectory into smaller line segments and clustering them based on similarity in terms of spatiotemporal properties. It can be useful when it is not necessary to analyze the entire trajectory. In contrast, whole-trajectory clustering is more commonly used for identifying patterns over a longer period of time. Lee et al. [31] introduced a novel partition-and-group framework for clustering sub-trajectories, which involves dividing trajectories into a collection of line segments and subsequently grouping similar line segments into a cluster [32]. Buchin et al. [33] presented a linear-time algorithm for calculating the most similar sub-trajectories under equal starting times, while for unequal starting times, they offered $(1+\varepsilon)$-approximation algorithms that were not as efficient. This approach employed a distance measure defined as the average Euclidean distance at corresponding times. Mao et al. [34] used a sliding window model to propose an online algorithm for clustering streaming trajectories. This model contained a micro-clustering that represented clustering trajectory line segments in a current window and a macro-clustering based on micro-clustering over a specified time interval. They defined two data structures for tracking the latest alterations of a cluster trajectory data stream in real-time. Buchin et al. [35] also examined the problem of identifying commuting patterns in trajectories by searching for comparable sub-trajectories. They proposed multiple approximation algorithms and demonstrated that computing and approximating the longest sub-trajectory cluster was as hard as the Max-Clique problem. Liu et al. [36] defined a method for determining the distance between sub-trajectories considering time, space, and direction. They proposed a density trajectory cluster algorithm in which the cluster radius is calculated by the density of the data distribution. Schreck et al. [37] suggested a complete visual-interactive monitoring and control framework for trajectory data analysis based upon an enhanced version of the Self-Organizing Maps algorithm. Their framework allowed users to observe clustering progression visually and manage the algorithm at any desired level. Yang et al. [38] designed a trajectory clustering algorithm based on spatio-temporal density analysis of data and defined two new metrics-NMAST density function and NT factor- in this algorithm. Binh Han et al. [39] proposed a systematic method for whole-trajectory clustering of mobile objects in road networks through their framework TRACEMOB. This includes a unique distance measure that captures the complex characteristics of trajectories, an algorithm to transform these trajectories into multidimensional data points in Euclidean space, and a validation method for assessing clustering quality. The validation approach evaluates the quality of clustering in both the transformed metric space and the original road network space, ensuring comprehensive analysis. Wang et al. [40] designed a grid-based whole-trajectory clustering algorithm that converts the trajectories in the road network space into corresponding grid sequences within grid space. Their approach could determine a series of clusters as well as some abnormal trajectories and GPS points.
@@
Anomaly detection in trajectories involves identifying patterns and behaviors that deviate significantly from the norm [41, 9, 42]. Though the deviations are big enough to qualify as statistical outliers, they are ignorable with respect to the inference on the overall clustering pattern. This is what we address in our work via the stable trajectory clustering algorithm, where the trajectories remain in their respective clusters to the extent possible despite minor anomalies in a few time intervals. However, in previous studies, stable clustering algorithms refer to methods that create and maintain clusters of data points to minimize changes in cluster membership over time. These algorithms ensure efficient communication and resource utilization while considering factors such as distance and connectivity. Shahwani et al. [43] proposed a novel clustering algorithm for Vehicular Ad hoc Networks (VANETs) that enhances cluster stability by utilizing vehicle trajectories, specifically incorporating the traffic regularity of buses, into the clustering process. The algorithm used the Affinity Propagation technique, which traditionally relies on distance metrics, but modified it to also consider the shared trajectories of vehicles. By doing so, the algorithm selects Cluster Heads (CHs), which are central nodes in VANETs responsible for coordinating communication within a cluster of vehicles. These CHs not only have minimal distance from other vehicles but also share longer common trajectories with them, leading to more stable clusters. Simulation results demonstrated that this trajectory-based approach significantly improved cluster lifetime and reduced the frequency of cluster head changes compared to conventional methods, thereby enhancing the overall efficiency and reliability of VANET communications. Tseng et al. [44] suggested a stable clustering algorithm called the Clustering Algorithm using the Traffic Regularity of Buses for VANETs in urban scenarios. The algorithm leveraged the fixed routes and regular schedules of buses to improve cluster stability. It considered vehicle mobility characteristics such as velocity, position, and direction, along with the spatial and temporal dependencies of traffic patterns in urban areas. CATRB used the concept of an equilateral triangle’s centroid, incenter, and circumcenter to select the most appropriate CH. This approach aimed to position the CH at the center of the cluster for efficient control and data transmission. These studies just focused on providing stable algorithms for data point clustering. While stable data point clustering is undoubtedly important, it does not fully address the complexities and dynamic nature of trajectory clustering. Our work aims to fill this gap by introducing and rigorously evaluating a stable algorithm designed for trajectory clustering.
III Preliminaries and Definitions

### v2

@@
Abstract.
Clustering algorithms fundamentally group data points by characteristics to identify patterns. Over the past two decades, researchers have extended these methods to analyze trajectories of humans, animals, and vehicles, studying their behavior and movement across applications. This paper presents whole-trajectory clustering and sub-trajectory clustering algorithms based on DBSCAN line segment clustering, which encompasses two key events: split and merge of line segments. The events are utilized to capture object movement history based on the average Euclidean distance between line segments. In this framework, whole-trajectory clustering considers entire entities’ trajectories, whereas sub-trajectory clustering employs a sliding window model to identify local similarity patterns. Many existing trajectory clustering algorithms respond to temporary anomalies in data by splitting trajectories, which often obscures otherwise consistent clustering patterns and leads to less reliable insights. To address this, we introduce the stable trajectory clustering algorithm, which leverages the mean absolute deviation concept to demonstrate that selective omission of transient deviations not only preserves the integrity of clusters but also improves their stability and interpretability. We evaluate all proposed algorithms on real trajectory datasets to illustrate their effectiveness and sensitivity to parameter variations.
@@ Abstract.
$c$s: Computing methodologies Cluster analysisccs: Theory of computation Computational geometryccs: Information systems Clusteringccs: Information systems Spatial-temporal systems
1. Introduction
@@
Trajectory clustering is one of the most important data analysis methods developed for analyzing moving objects like humans, animals, and vehicles in various fields and applications. It can be used to identify common behaviors and recognize similar movement patterns, as well as to detect irregular or anomalous movements in different types of trajectory data. Among the use cases of trajectory clustering, we can mention user lifestyle profiling to capture behavioral similarities beyond geographic location (63), exploring urban travel patterns that reflect the regularity of mobility patterns, which is an important factor in evaluating the optimality of city spatial structure and road network connectivity (7), classifying animal behaviors (foraging, resting, walking) using high-frequency tracking data to study their environmental interactions (6) and clustering animals by species based on their daily movement patterns, which is complicated by the erratic nature of these movements (2). Trajectory clustering is also used in air traffic categorization (3; 5), traffic network identification (4), and weather prediction (8).
@@
Anomalies frequently occur in trajectory clustering due to irregularities or deviations in movement patterns. To address this, numerous algorithms have been developed, each designed to effectively detect and analyze these anomalies (20; 46). Missing, however, is an algorithm that can ignore insignificant anomalies in trajectories to reveal clustering patterns that are otherwise stable. Current methods usually handle these deviations in two ways, but both have limitations. The first strategy is to preprocess the raw trajectory data to remove small deviations. This is not ideal because it might delete important information and change the true nature of the trajectory. A more common strategy is to split a trajectory from its cluster as soon as a deviation appears. It often creates fragmented clusters, which hide the main movement pattern and make the results hard to understand. In recent years, researchers have focused on stable algorithms for clustering data points that select the optimal number of clusters (18). In other words, current stable clustering algorithms try to split and merge data points correctly in a way that is robust against random oscillations. Inspired by advances in stable clustering of data points, we propose a stable algorithm for trajectory clustering that can be applied on existing trajectory clustering methods. In this paper, we apply the stable algorithm to a trajectory clustering approach based on line segment clustering, which naturally produces split and merge events over time. Earlier studies also employed line segment clustering, but without incorporating temporal labels or explicitly defining split and merge operations in trajectory clustering (16; 17). The core contribution of our paper is the stable trajectory clustering algorithm itself, which efficiently improves cluster quality by determining when outlier trajectories can be assigned to their respective clusters despite insignificant anomalies. It addresses a prominent issue in real trajectory data that previous algorithms have not addressed.
@@
Building upon existing trajectory clustering algorithms, we propose algorithms for both whole-trajectory clustering and sub-trajectory clustering. A trajectory can be described as a piecewise linear function, where each segment represents the movement of an object over a time interval. Consequently, for trajectory clustering, a set of line segments exists for each time interval, and these sets need to be clustered. For this purpose, Density-Based Spatial Clustering of Applications with Noise (DBSCAN) line segment clustering defines the split and merge procedures of line segments. One key advantage of this approach is that the results from clustering in each time interval can be reused in the subsequent interval, as line segments often retain their previous cluster assignment. The clustering process relies on the continuous average Euclidean distance as the similarity measure between line segments. Therefore, a data analyst can identify the behaviors of moving objects across all time intervals and determine which objects move together and belong to the same cluster in each interval. Trajectories whose line segments belong to the same cluster and show similar behaviors in each time interval are eventually grouped into a single whole-trajectory cluster. In addition to whole-trajectory clustering, the proposed framework includes a sub-trajectory clustering algorithm that focuses on specific time intervals to identify similar sub-trajectories. It employs a sliding window model with two parameters to capture temporal patterns and discover recurrent sub-sequences. We then present an innovative definition and algorithm for the stable trajectory clustering algorithm, which is a first-of-its-kind advancement. The algorithm ignores minor and insignificant deviations of trajectories (outlier trajectories) that temporarily leave their clusters and then return to them. The approach reflects the true movement patterns more accurately, which makes the clusters more meaningful and easier to interpret. Comparable situations occur in real-world movement patterns where short departures do not change the general movement pattern. For instance, in urban traffic analysis, a vehicle making a brief stop at a traffic light or a pedestrian crossing should still be classified within the main traffic cluster. Furthermore, a person briefly deviating from a common trajectory to look at a shop window should not be marked as an outlier in pedestrian movement studies. A related example can be seen in animal movement studies, where a migrating bird or herd of animals temporarily deviates from its migration trajectory to find food or rest, but then returns to their original trajectory afterward. In this algorithm, we employ the mean absolute deviation concept to assess how much each outlier trajectory deviates from its assigned cluster, which enables us to distinguish significant outliers from temporary anomalies. In the final stage, all algorithms are applied to two real-world trajectory datasets to analyze parameter sensitivity and assess the effect on the number of clusters. Three evaluation criteria are then employed to measure clustering performance before and after running the stable trajectory clustering algorithm: the Silhouette score evaluates the quality of the internal cluster, and Normalized Mutual Information (NMI), together with the Adjusted Rand Index (ARI), provides external benchmarks for comparison with the traditional and deep learning methods. The results show that the stable trajectory clustering algorithm efficiently manages insignificant anomalies to improve the quality of clusters and achieve better clustering accuracy compared to other approaches. In summary, our contributions are illustrated as a graphical abstract in Fig. 1.
@@
Figure 1. Graphical abstract of the proposed approaches.
2. Related Work
@@
In the last two decades, many data point clustering algorithms have been discussed in various fields and applications such as wireless sensor networks (32), flight anomaly detection (33), and image processing (34). These algorithms try to group similar data points based on their similarities or dissimilarities (23). There are several types of methods to cluster data points, such as partitioning methods (25; 24; 26), hierarchical methods (28; 29), grid-based methods (30; 31), and density-based methods (22; 27). Some of these methods are used for clustering data points as well as for trajectory clustering. For example, DBSCAN (22) is a density-based algorithm that is used for clustering both data points and trajectories. It can identify the shape and size of data, can handle noise and outliers efficiently, and does not require the number of clusters as a user-defined input. However, DBSCAN cannot identify clusters with different densities and is sensitive to the choice of user input parameters for clustering. Based on the algorithms, researchers then started developing new methods for clustering continuous movement objects, known as trajectory clustering.
@@
Trajectory data are recorded in different formats by various devices such as Global Positioning System (GPS) trackers, Radio Frequency Identification (RFID) systems, and smartphones. Numerous algorithms have been designed to cluster these different types of trajectory data. They also contribute to improving research, innovation, and theoretical understanding in various fields and applications. Trajectory clustering can be divided into point-based, sub-trajectory-based, and whole-trajectory-based methods. Point-based trajectory clustering is a method of clustering trajectories by comparing specific locations (points) along each trajectory. Instead of looking at the entire trajectory, it focuses on the individual points where the object was at different times and groups similar ones together. Li et al. (12) offered a group discovery framework to facilitate the efficient online discovery of moving objects traveling collectively. Their framework employed a sampling-independent approach while also naturally supporting the utilization of approximate trajectories. Jensen et al. (11) investigated effectively managing a collection of constantly moving data points in a two-dimensional Euclidean space to maintain a cluster of dynamic data points. Their proposed approach can cluster moving objects incrementally. Tang et al. (10) introduced the smart-and-close discovery algorithm, which efficiently generates companions from trajectory data. In their approach, the objective of the proposed traveling buddy model was to improve the clustering and intersection processes involved in discovering companions. Li et al. (9) recommended utilizing cluster analysis for mobile objects and introduced the notion of moving micro-clusters, which can effectively identify patterns in the mobility of objects even in massive datasets. Gao et al. (48) addressed the challenges of clustering real-time trajectory data generated by GPS-equipped devices. They proposed a novel framework that dynamically models trajectory streams using a self-adaptive k-segment method, considering both spatial and temporal dimensions. This framework supported distributed indexing and processing to handle large-scale data efficiently. It utilized incremental DBSCAN clustering for real-time processing and integrated a two-layered distributed index to enhance scalability. The study’s key contributions included effective trajectory stream modeling, efficient and incremental clustering algorithms, and scalable distributed processing techniques.
@@
Sub-trajectory clustering involves breaking down a trajectory into smaller line segments and clustering them based on similarity in terms of spatiotemporal properties. It can be useful when it is not necessary to analyze the entire trajectory. In contrast, whole-trajectory clustering is more commonly used for identifying patterns over a longer period of time. Lee et al. (16) introduced a novel partition-and-group framework for clustering sub-trajectories, which involves dividing trajectories into a collection of line segments and subsequently grouping similar line segments into a cluster (13). However, their framework does not consider temporal labels or the evolution of clusters across time, and therefore does not capture split and merge dynamics in trajectory clustering. Yu et al. (17) suggested a density-based clustering algorithm for trajectory data streams. They designed an algorithm to handle moving object trajectory data as an incremental line segment stream, using a TC-Tree structure to update clusters in real-time upon the arrival of each new segment. This is distinct from our interval-based approach, which analyzes the evolution of entire clusters between discrete time intervals. Da et al. (38) dealt with the problem of identifying sub-trajectory clusters at every time interval. To this aim, they defined a structure called a micro-group to take and maintain small groups of moving objects at each time window and presented an algorithm for maintaining each micro-group incrementally to discover patterns. Thus, they suggested a density-based sub-trajectory cluster identification approach by merging micro-groups. Buchin et al. (14) presented a linear-time algorithm for calculating the most similar sub-trajectories under equal starting times, while for unequal starting times, they offered $(1+\varepsilon)$-approximation algorithms that were not as efficient. This approach employed a distance measure defined as the average Euclidean distance at corresponding times. Mao et al. (36) used a sliding window model to propose an online algorithm for clustering streaming trajectories. This model contained a micro-clustering that represented clustering trajectory line segments in a current window and a macro-clustering based on micro-clustering over a specified time interval. They defined two data structures for tracking the latest alterations of a cluster trajectory data stream in real-time. Buchin et al. (15) also examined the problem of identifying commuting patterns in trajectories by searching for comparable sub-trajectories. They proposed multiple approximation algorithms and demonstrated that computing and approximating the longest sub-trajectory cluster was as hard as the Max-Clique problem. Liu et al. (35) defined a method for determining the distance between sub-trajectories considering time, space, and direction. They proposed a density trajectory cluster algorithm in which the cluster radius is calculated by the density of the data distribution. Schreck et al. (37) suggested a complete visual-interactive monitoring and control framework for trajectory data analysis based upon an enhanced version of the Self-Organizing Maps algorithm. Their framework allowed users to observe clustering progression visually and manage the algorithm at any desired level. Yang et al. (1) designed a trajectory clustering algorithm based on spatio-temporal density analysis of data and defined two new metrics-NMAST density function and NT factor- in this algorithm. Binh Han et al. (45) proposed a systematic method for whole-trajectory clustering of mobile objects in road networks through their framework TRACEMOB. This includes a unique distance measure that captures the complex characteristics of trajectories, an algorithm to transform these trajectories into multidimensional data points in Euclidean space, and a validation method for assessing clustering quality. The validation approach evaluates the quality of clustering in both the transformed metric space and the original road network space, ensuring comprehensive analysis. Wang et al. (21) designed a grid-based whole-trajectory clustering algorithm that converts the trajectories in the road network space into corresponding grid sequences within grid space. Their approach could determine a series of clusters as well as some abnormal trajectories and GPS points.
In addition to the aforementioned methods, some studies have used deep learning for trajectory clustering. Yao et al. (59) proposed a deep learning based approach to improve trajectory clustering. They used a sequence-to-sequence autoencoder to turn each trajectory into a short vector, which represents its overall movement pattern. Their method then clusters these simple vectors instead of the complex raw trajectories. An advantage of this method is that it allows the model to group trajectories with similar shapes, even if they aren’t perfectly aligned in time or have different lengths. Furthermore, Wang et al.(60) introduced an end-to-end framework that learns how to represent trajectories and cluster them at the same time. Their method, called DTC, also relies on a sequence-to-sequence autoencoder that turns each trajectory into a vector. In other words, the model is trained to do two things at once: learn a good vector for each trajectory, and also encourage the vectors of similar trajectories to be closer together into clusters. The main new idea is that the model doesn’t work in two steps. Instead, it learns the representations and groups them into clusters at the same time, which helps improve the final result. Building on this end-to-end approach, Wang et al.(61) proposed a more advanced framework for spatiotemporal trajectory clustering called DSTC. Their work argues that accurate clustering requires a more sophisticated representation of both space and time. To achieve this, for space, it replaces the rigid grid with a flexible, density-based approach that adapts to how data is distributed. For time, it introduces polar coordinates to correctly represent its cyclical nature. Second, it explicitly models the cyclical nature of time using polar coordinates. These improvements lead to state-of-the-art clustering accuracy. However, achieving this performance requires a much more complex model that is heavily reliant on very large datasets.
@@
Although clustering techniques have advanced, many methods still find it hard to deal with irregularities in trajectories. Anomaly detection in trajectories involves identifying patterns and behaviors that deviate significantly from the norm (19; 46; 50; 62). Though the deviations are big enough to qualify as statistical outliers, they are ignorable with respect to the inference on the overall clustering pattern. This is what we address in our work via the stable trajectory clustering algorithm, where the trajectories remain in their respective clusters to the extent possible despite minor anomalies in a few time intervals. However, in previous studies, stable clustering algorithms refer to methods that create and maintain clusters of data points to minimize changes in cluster membership over time. These algorithms ensure efficient communication and resource utilization while considering factors such as distance and connectivity. Shahwani et al. (47) proposed a novel clustering algorithm for Vehicular Ad hoc Networks (VANETs) that enhances cluster stability by utilizing vehicle trajectories, specifically incorporating the traffic regularity of buses, into the clustering process. The algorithm used the Affinity Propagation technique, which traditionally relies on distance metrics, but modified it to also consider the shared trajectories of vehicles. By doing so, the algorithm selects Cluster Heads (CHs), which are central nodes in VANETs responsible for coordinating communication within a cluster of vehicles. These CHs not only have minimal distance from other vehicles but also share longer common trajectories with them, leading to more stable clusters. Simulation results demonstrated that this trajectory-based approach significantly improved cluster lifetime and reduced the frequency of cluster head changes compared to conventional methods, thereby enhancing the overall efficiency and reliability of VANET communications. Tseng et al. (54) suggested a stable clustering algorithm called the Clustering Algorithm using the Traffic Regularity of Buses for VANETs in urban scenarios. The algorithm leveraged the fixed routes and regular schedules of buses to improve cluster stability. It considered vehicle mobility characteristics such as velocity, position, and direction, along with the spatial and temporal dependencies of traffic patterns in urban areas. CATRB used the concept of an equilateral triangle’s centroid, incenter, and circumcenter to select the most appropriate CH. This approach aimed to position the CH at the center of the cluster for efficient control and data transmission. These studies just focused on providing stable algorithms for data point clustering. While stable data point clustering is undoubtedly important, it does not fully address the complexities and dynamic nature of trajectory clustering. Our work aims to fill this gap by introducing and rigorously evaluating a stable algorithm designed for trajectory clustering.
3. Preliminaries and Definitions